OpenStudy (anonymous):

Prove the identity by changing the left side of the equation to show equivalency to the right side. (1+cos2ø)/(sin2ø) = cotø

OpenStudy (solomonzelman):

\[(1+Cos^2x)(Sin^2x)=Cot^2x\]

OpenStudy (solomonzelman):

duagles take over, i got replied....

OpenStudy (anonymous):

I'm stumped. Perhaps this is cos(2x) rather than (cosx)^2? etc? sin(2x) = sin(x+x) = sinx cosx + sinx cosx =2 sin x cos x

OpenStudy (schrodingers_cat):